1.     Producers in three countries have formed a cartel to sell a highly-priced product. Suppose the world demand for the product is consistent with the equation:

P=80-0.005QT

 

NB: QT is the total cartel sales.

The producers have managed to maximise profit at a price of R70 per unit. Suppose the members have the following total cost functions;

Member 1: TC=7382 + 10Q + 0.05Q2

Member 2: TC= 9374 + 20Q + 0.02Q2

Member 3: TC=7432 + 30Q + 0.03Q2

 

Tasks: Derive:

a)     Marginal revenue

b)    Output of member 1

c)     Output of member 2

d)    Output of member 3

e)    Total cartel output.


1
Expert's answer
2021-09-04T15:17:01-0400

a)     Marginal revenue

We know that the producers have managed to maximize profit at a price of R70 per unit.

"TR = P\times Q\\ 

TR = 70Q"

TR is the same for all the producers and the world demand for the product is consistent with :

"P=80-0.005Q_T\\MR = \frac{dTR}{dQ}"

Now, the world demand for the product "P=80-0.005Q_T"

"TR = (80-0.005Q_T) Q_T\\

 TR = 80Q_T - 0.005Q_T^2\\

 MR = 80 - 0.01 Q_T"


b)    Output of member 1

"member \space 1:TC=7382 + 10Q_1 + 0.05Q_1^2\\

 MC = \frac{dTC}{dQ}\\
 MC = 10 + 0.1Q_1"

 The producer will produce that output where MC=MR

"10 + 0.1Q_1 = 70\\

 0.1 Q_1 = 70 - 10\\

 0.1 Q_1 = 60 \\

 Q_1 = 600 \space units"


c)     Output of member 2

"member \space 2:TC=9374 + 20Q_2 + 0.02Q_2^2\\

 MC = \frac{dTC}{dQ}\\
 MC = 20 + 0.04Q_2" 

The producer will produce that output where MC=MR

"20 + 0.04Q_2 = 70\\

 0.04 Q_2 = 70 - 20\\

 0.04 Q_2 = 50 \\

 Q_2 = 1250 \space units"


d)    Output of member 3

"member \space 3:TC=7432 + 30Q_3 + 0.03Q_3^2\\

 MC = \frac{dTC}{dQ}\\
 MC = 30 + 0.06Q_3" 

The producer will produce that output where MC=MR

"30 + 0.06Q_3 = 70\\

 0.06 Q_3 = 70 - 30\\

 0.06 Q_3 = 40 \\

 Q_3 = 666.67 \space units\\ Q_3 = 667 \space units"


e)    Total cartel output.

Total cartel output = Output of all the three producers members

"= Q_1 + Q_2 + Q_3\\

 = 600 + 1250 + 667\\

 = 2517 \space units"

Total cartel Output = 2517 units.